For the diprotic weak acid H2A, Ka1 = 3.5 × 10-6 and Ka2 = 6.2 × 10-9. What is the pH of a 0.0800 M solution of H2A? What are the equilibrium concentrations of H2A and A2– in this solution?
Answer #1

apply

H2A <-> H+ +HA-

HA- <-> H+ + A-2

Ka1 = [H+][HA-]/[H2A]

3.5*10^-6 = (x*x)/(0.08-x)

x = 5.27*10^-4

then

Ka2 =  [H+][A-2] / [HA-]

Ka2 = (y*y)/(x-y)

6.2*10^-9 = y*y/(5.27*10^-4 -y)

y = 1.8*10^-6

then

Htotal = x+y = 5.27*10^-4+1.8*10^-6 = 0.0005288

pH= -log(0.0005288) = 3.276

H2A = M-x = 0.0800 -5.27*10^-4 = 0.079473 M

A-2 = y = 1.8*10^-6 M